Contribution Margin by Segment

The following data are for Moss Creek Apparel:

East West
Sales volume (units):
    Product XX 7,500 7,500
    Product YY 2,500 4,100
Sales price:
    Product XX $23 $22
    Product YY $27 $28
Variable cost per unit:
    Product XX $14 $14
    Product YY $16 $16

a. Determine the contribution margin for Product YY.
$

b. Determine the contribution margin for the West Region.
$
Answer #1

a. Determine the contribution margin for Product YY

= Contribution margin of Product YY (East Region) + Contribution margin of Product YY (West region)

Contribution margin of product YY (East Region) = units * (Selling price - variable cost)  

= 2,500 * ($27 - $16)

= $27,500

Contribution margin of product YY (East region) = 4,100 * ($28 - $16)

= $49,200

Contribution margin of Product YY = $27,500 + $49,200

= $76,700

b. Determine the contribution margin for the West region

= Contribution margin of Product XX (West region) + Contribution margin of product YY(West region)

Contribution margin of product XX (West region) = 7,500 * ($22 - $14)

= $60,000

Contribution margin of product YY(West region) = 4,100 * ($28 - $16)

= $49,200

Contribution margin for the West region = $60,000 +$49,200

= $109,200